# Dynamic Routing Between Experts: A Data-Efficient Approach to Continual Learning in Vision-Language Models

### v2

@@ Results
Table 1: Comparison of continual learning methods on specialized and benchmark tasks using InternVL2-2B.
As shown in Table 1, InternVL2-2B demonstrates strong zero-shot performance on benchmark tasks, but performs relatively poorly on specialized tasks, highlighting the need for task-specific adaptation. Traditional sequential fine-tuning improves performance on the most recent task but suffers from severe CF, leading to substantial degradation on earlier tasks and benchmarks, consistent with prior findings (Goodfellow et al. 2013; Wang et al. 2024b).
To address this limitation, we incorporate experience replay by retaining 20% of data from previous tasks. This approach improves stability over naive sequential fine-tuning but still results in performance drops—particularly on SNLI-VE—and requires storing historical data, which increases computational overhead as tasks accumulate. Both sequential fine-tuning strategies show significant degradation on benchmark tasks, highlighting the model’s inability to retain general-purpose knowledge.
Next, we explore an efficient modular approach: model merging. We evaluate two strategies: model weight averaging and task vector arithmetic. These methods outperform traditional sequential fine-tuning while eliminating the dependency on historical data, requiring only the adaptation of the base model to new tasks. These methods require only task-specific adaptation of the base model and effectively preserve the VLM’s general-purpose capabilities, with performance closely matching the zero-shot baseline. However, despite these advantages, a notable performance gap remains when compared to MTL on specialized tasks.
In contrast, our routing-based approach achieves performance on specialized tasks that is comparable to or better than MTL, while preserving the model’s original benchmark performance. For example, on MMIMDb, routing even surpasses MTL, possibly due to reduced negative task interference (Wang, Lipton, and Tsvetkov 2020). Furthermore, while MTL suffers a substantial drop on benchmark tasks—indicating loss of general-purpose capabilities—routing maintains performance near the zero-shot baseline. This demonstrates that routing not only mitigates catastrophic forgetting, but also balances specialization and generalization without the computational and data burdens of multi-task training.
Method Specialized Tasks Benchmark Tasks
COCO SNLI HM MMIMDb MMB ChartQA DocVQA
Zero-shot 0.889 70.2 65.4 54.5 85.6 72.6 89.7
Merging: Task Arith. 1.272 71.3 68.7 66.9 86.0 70.6 87.2
Routing 1.350 83.3 76.8 67.4 86.3 73.1 89.3
@@
Table 2: Comparison of top-performing continual learning approaches for InternVL2-8B on specialized and benchmark tasks.
InternVL2-8B Results. Table 2 extends our evaluation to the larger InternVL2-8B model, focusing on the top-performing methods from the 2B setting: task arithmetic, routing, and MTL. At this scale, routing matches MTL in performance, further closing the gap observed with the smaller model, while maintaining benchmark performance without degradation. Notably, routing not only preserves general-purpose capabilities but also surpasses the zero-shot baseline—highlighting that its effectiveness scales with model size. These results suggest routing becomes increasingly advantageous as model capacity grows.
Building on findings from Yadav et al. (2024), which show that model merging improves with scale, we also conducted model merging experiments using the 8B model. Compared to the 2B setting, merging indeed yielded better performance. However, it still lagged behind routing and MTL. One key limitation of model merging is its tendency to disproportionately favor tasks with more training data, allowing their influence to dominate the merged weights. For example, in our experiments, captioning performance was retained significantly better than that on SNLI-VE (SNLI) or Hateful Memes (HM), which had fewer examples.
Routing Matches or Outperforms Specialized Models
@@
Routing 1.350 83.3 76.8 67.4 86.3 73.1 89.4
Table 3 compares routing against both zero-shot and task-specific LoRA fine-tuned models across specialized for InternVL-2 8B model. While LoRA models offer strong performance on the task they were fine-tuned for (e.g., LoRA-SNLI excels on SNLI), they often exhibit reduced performance on unrelated domains, revealing limited generalization.
In contrast, the routing model matches or surpasses each specialized LoRA variant on its corresponding task—achieving better scores on MMIMDb and Hateful Memes—and performs competitively on others. Notably, routing maintains strong results across benchmark tasks such as MMBench and ChartQA, rivaling or slightly exceeding even the zero-shot and LoRA baselines. This demonstrates its capacity to adaptively activate relevant modules while preserving the base model’s general-purpose strengths.
By learning token-level gates that modulate each LoRA module’s contribution, routing avoids the interference typically caused by multi-task fine-tuning and mitigates catastrophic forgetting. Its selective activation mechanism enables it to leverage task-specific expertise without compromising robustness across unrelated domains—delivering the benefits of specialization and generalization in a single model.
Ablation Study
In this section we perform different ablations on routing based approach to demonstrate it’s robustness and computational efficiency compared to other approaches.
Adding more LoRA modules doesn’t degrade performance
In this section, we evaluate the scalability and robustness of the routing mechanism as more LoRA modules are incrementally added to the model. Our goal is to test whether introducing additional experts causes performance degradation or interference among tasks.
We begin by routing between two task-specific experts, SNLI and COCO. As shown in Table 4, this configuration yields strong performance on both SNLI and COCO while having nearly zero-shot performance on unrelated tasks like Hateful Memes and MMIMDb.
@@
Adding a LoRA module for the Hateful Memes task leads to a substantial gain in that task’s performance, demonstrating the model’s capacity to correctly route samples to the appropriate expert. Extending this setup by including the MMIMDb LoRA module results in significant improvement on MMIMDb, while maintaining previously attained performance levels across other tasks.
To test the system’s robustness to unrelated modules, we add a LoRA module trained on ChartGEMMA (Masry et al. 2024), a dataset unrelated to the core tasks. As seen in Table 4, the inclusion of this unrelated module does not degrade performance, reinforcing the routing system’s resilience and scalability, even in the presence of unrelated modules.
Finally, we introduce a LoRA module for XNLI (Conneau et al. 2018), a text-to-text entailment task closely related to SNLI, which specializes in text-to-image entailment task. Interestingly, adding the XNLI module not only boosts performance on image-text tasks but also shows the model’s ability to leverage cross-modal transfer, where improvements on text-based tasks benefit performance on related image-based tasks.
@@
orca-math (text) multi-math (image+text)
@@
Table 5: Cross-modal transfer experiments using routing with modality-specific experts on the MGSM dataset.
Routing Enables Cross-modal Transfer
Building on our findings from the XNLI and SNLI experiments, we further examine routing’s effectiveness in enabling cross-modal transfer within VLMs. Specifically, we evaluate this on the multi-lingual math dataset MGSM (Shi et al. 2022), focusing on zero-shot performance.
As shown in Table 5, we employ two LoRA modules: Orca-Math, trained on text-based mathematical reasoning, and Multi-Math, trained on both image and text modalities. Both modules improve performance on MGSM, with Orca-Math outperforming Multi-Math, likely due to its alignment with the task structure of MGSM.
Crucially, routing achieves the strongest performance across both languages by dynamically integrating knowledge from multiple experts, surpassing individual LoRA modules. This highlights its effectiveness in cross-modal knowledge transfer, wherein model is able to use information from both text based orca-math expert as well as image and text based mutli-math expert allowing the model to leverage complementary information from diverse modalities. These results demonstrate that routing serves as a robust mechanism for improving multi-modal performance by effectively leveraging knowledge from unimodal tasks.
Impact of Model Size
In this section, we examine the effect of model size on the routing approach by comparing the performance of specialized models with that of a routing model. The routing model integrates SNLI-VE, COCO, Hateful Memes, and MMIMDb as LoRA modules within the routing pool. A smaller performance drop in the routing model, relative to specialized models, indicates better robustness.
Figure 2: Performance drop in comparison to specialized model for 2B and 8B models.
In Figure 2, the 8B model exhibits superior robustness, with minimal performance drops across all tasks compared to the 2B model. Notably, the 8B routing model even outperforms the specialized models on tasks such as COCO, Hateful Memes, and MMIMDb. In contrast, the 2B model experiences a more significant performance drop across all tasks, with SNLI-VE showing the largest gap.
These results indicate that larger models, such as the 8B variant, can effectively manage multiple tasks with minimal trade-offs, likely due to their increased capacity and improved generalization capabilities. This behavior is consistent with findings from prior research on model merging, as noted by Yadav et al. (2024).
Computational Efficiency Analysis
We analyze the computational advantages of our routing-based LoRA approach compared to standard continual learning methods. Table 6 reports training time and data requirements for InternVL2-2B across four specialized tasks.
Method Training Time (hrs) Data Requirements
First Task All Tasks First Task All Tasks
Seq. FT 8 11.0
Seq. FT + ER 8 12.8 2.2×
MTL 8 15.6
Routing 4 8.4
Table 6: Comparison of computational efficiency. Training time in GPU hours on 32×A10 GPUs; data usage is shown relative to the first task.
As shown in Table 6, our routing-based approach is substantially more efficient in both training time and data usage. This efficiency arises from two key design choices: (i) the use of lightweight LoRA adapters, and (ii) training exclusively on the current task’s data, without revisiting prior datasets or retraining earlier components. While multi-task learning (MTL) can match routing in specialized task performance, it incurs significantly higher compute and data costs—approximately 2× training time and 4× data usage—as it requires joint access to all task data and, as shown in Table 1, suffers from degraded performance on benchmark tasks due to CF on general purpose knowledge. Experience replay (ER), although more efficient than MTL, still faces growing complexity as the number of tasks increases. We demonstrate that routing framework strikes a practical balance by maintaining strong performance, limiting computational overhead, and scaling easily to an increasing number of tasks—making it well-suited for real-world continual learning scenarios.
Discussion
Relationship to Mixture of Experts
While our routing approach shares architectural similarities with Mixture of Experts (MoE) systems like MoE-LLaVA (Chen et al. 2024a), it fundamentally differs in both purpose and implementation. It’s important to clarify these distinctions as they represent different paradigms in modular machine learning.
Traditional MoE approaches are designed to improve model capacity and performance through joint training of specialized experts within a unified architecture. During both training and inference, a gating network distributes computation across experts based on input characteristics. This requires simultaneous access to all training data and joint optimization of all experts.
In contrast, our routing-based continual learning approach:
  1. 1.
    Sequential vs. Joint Training: Our experts (LoRA modules) are trained independently and sequentially, with each expert focusing on a specific task without access to other tasks’ data. Traditional MoEs require joint training of all experts from the beginning.
  2. 2.
@@ Discussion
    Architectural Separation: Our approach explicitly decouples expert learning from routing. We first train task-specific LoRA modules independently, and then learn lightweight routing vectors while keeping the LoRA parameters frozen. This modular design supports continual adaptation, unlike traditional MoE systems where expert structure and routing are jointly optimized from the get-go on all tasks.
  3. 3.
    Data Efficiency: Our approach requires only new task data, eliminating the need to retain or synthesize previous task data—a key requirement for practical continual learning in data-constrained environments.
Overall, unlike traditional MoEs that route among fully joint-trained large experts, our approach routes among lightweight, task-specialized parameter-efficient modules optimized for scalable continual adaptation in large vision-language models. This makes our method particularly well-suited for settings where full retraining of large experts is computationally prohibitive or data reuse is restricted.
Conclusion
In this work, we address catastrophic forgetting in Vision-Language Models (VLMs) by introducing a routing-based continual learning approach that enables seamless integration of new task-specific experts while preserving foundational capabilities. Unlike multi-task learning and experience replay, which incur $\mathcal{O}(n)$ growth in data and computation, our method maintains constant $\mathcal{O}(1)$ resource requirements per task and eliminates the need to revisit prior data, making it highly efficient. Through extensive experiments with InternVL2 models (2B and 8B), we show that routing achieves performance comparable to multi-task learning—the theoretical upper bound—on specialized tasks, while preserving general capabilities on benchmarks such as ChartQA, DocVQA, and MMBench. Routing also enables cross-modal transfer, allowing improvements in one modality (e.g., vision) to enhance performance in another (e.g., language), a property not exhibited by existing continual learning methods. Furthermore, we find that larger models exhibit increased resistance to forgetting. In future work, we aim to extend this framework to support cross-lingual transfer in VLMs.

### v3

@@ 4 Experiments
Routing (Ours) 1.305 80.3 71.2 63.2 78.8 59.2 84.3
8B Zero-shot 0.889 70.2 65.4 54.5 85.6 72.6 89.7
Seq-FT 1.258 65.4 52.1 64.8 71.2 48.5 74.3
Arith. Merge 1.272 71.3 68.7 66.9 86.0 70.6 87.2
MR-LoRA 1.351 81.0 73.4 66.7 - - -
@@
Routing (Ours) 1.350 83.3 76.8 67.4 86.3 73.1 89.3
4.1 Datasets and Evaluation Metrics
To evaluate the stability-plasticity trade-off, we categorize our evaluation into Specialization Tasks and General Benchmarks.
Specialization Tasks.
We employ a diverse set of multimodal (image+text) tasks covering: (1) COCO-Caption (COCO) Lin et al. (2014) for image captioning (CIDEr); (2) SNLI-VE Xie et al. (2019) for visual entailment (Accuracy); (3) Hateful Memes (HM) Kiela et al. (2020) for toxicity detection (Accuracy); and (4) MMIMDb Arevalo et al. (2017) for movie genre classification (Precision).
General Benchmarks.
We monitor retention of foundational knowledge using MMBench (MMB) Liu et al. (2024b) for multi-dimensional reasoning, ChartQA Masry et al. (2022) for visual data interpretation, and DocVQA Mathew et al. (2021) for document understanding.
MLLM-CL benchmark.
Zhao et al. (2025) This benchmark evaluates the model’s ability to sequentially acquire knowledge across highly specialized and disparate domains, including Remote Sensing (RS), Medical imaging (Med), Art and Design (AD), Science (Sci), and Finance (Fin).
Cross-Modal Transfer.
To investigate the cross-modal transfer capabilities of the proposed method, we train two distinct experts: a text-only math expert (Orca-Math Mitra et al. (2024)) and a multimodal math expert (Multi-math Peng et al. (2024)). We then evaluate zero-shot performance on the MGSM framework Shi et al. (2022) in both English and Chinese.
4.2 Training Hyperparameters
Baseline Configurations.
For Sequential Finetuning (Seq-FT), we use a learning rate of $1\times 10^{-5}$ for three epochs on each task. For Multi-Task Learning (MTL), all tasks are co-trained for three epochs. All experiments are conducted on 32 A10 GPUs with a global batch size of 512. Arithmetic Merging (Arith. Merge) Ilharco et al. (2022) computes a weighted average finetuned weights to create a single model.
Modular Routing Training.
We follow the two-stage training strategy as discussed in Section 3:
  1. 1.
    Expert Specialization: Task-specific LoRA modules are trained ($lr=4\times 10^{-5}$, cosine decay) within the LLM backbone.
  2. 2.
    Gate Optimization: We freeze the LoRA weights and train the routing vector $v$ for 100 steps on the task data with batch size of 32 to map the expert to its specific activation manifold.
Task Sequence.
We evaluate continual learning using the sequence: $\text{COCO}\rightarrow\text{SNLI}\rightarrow\text{HM}\rightarrow\text{MMIMDb}$. During inference, we employ top-$k$ routing with $k=1$, prioritizing the most relevant expert for each token to maintain computational efficiency. We follow the suggested tasks sequence for MLLM-CL benchmark Zhao et al. (2025).
5 Results and Analysis
5.1 Performance Comparison and Stability
As demonstrated in Table 1, the proposed routing method (Routing) consistently achieves the optimal balance between task-specific specialization and the preservation of foundational capabilities.
Routing vs. Multi-Task Learning.
Notably, our routing approach achieves performance comparably with the MTL model. On MMIMDb, routing surpasses MTL in both the 2B (63.2 vs. 63.0) and 8B (67.4 vs. 65.4) variants. This suggests that routing effectively mitigates negative task interference Wang et al. (2020). While MTL forces a shared weight space to accommodate contrasting data distributions, our approach isolates task-specific features, preventing destructive gradient interference.
Furthermore, routing demonstrates superior stability. MTL exhibits performance degradation on MMBench and DocVQA compared to the zero-shot baseline, indicating a forgetting of foundational knowledge. In contrast, routing preserves these scores near zero-shot levels.
Routing vs. Arith. Merge.
Our experiments reveal a persistent performance gap between merging and routing on specialized tasks. While arithmetic merging effectively preserves general-purpose capabilities, it fails to fully recover the specialized performance of individual LoRA modules. For example, in the 8B setting, routing outperforms merging by $12.0\%$ on the SNLI task (83.3 vs. 71.3).
Routing vs. Seq-FT.
Sequential FT exhibits the most severe catastrophic forgetting, with performance on general benchmarks like MMBench dropping sharply (from 79.1 to 67.1 in the 2B model). While merging and MTL offer varying degrees of stability, they either sacrifice peak task performance or require simultaneous access to all training data. Routing achieves a “best-of-both-worlds” outcome: it doesn’t require access to past training dataset while matching or exceeding the stability of global multi-task training.
@@ 5.1 Performance Comparison and Stability
Routing (Ours) 1.350 83.3 76.8 67.4 86.3 73.1 89.4
Comparison with MR-LoRA.
Our framework offers several key advantages over MR-LoRA Zhao et al. (2025). Although MR-LoRA achieves strong task specialization, it requires a dual-inference pass, which increases computational overhead. In contrast, our routing mechanism is context-aware and completes inference in a single pass. Moreover, our approach benefits from cross-task synergy: by operating at the token level, the router can dynamically leverage specialized knowledge from previously learned tasks to assist in new domains. This capability enables routing to outperform MR-LoRA.
5.2 Specialization vs. Generalization
A key question in modular architectures is whether a single routed model can match the performance of individual, fully specialized models. Table 2 compares our framework against single-task LoRA experts.
The results indicate that routing does not merely match the performance of its experts; it often outperforms the individual LoRA variants from which it is composed. For example, the routed model achieves a score of 76.8 on Hateful Memes, surpassing the 76.5 achieved by the standalone LoRA-HM model. This suggests that the routing mechanism can perform "cross-expert utilization," where tokens within a single input benefit from the specialized features of multiple expert modules.
Table 3: Performance comparison on the MLLM-CL benchmark using LLaVA. Comparison is restricted to purely incremental approaches. Results are rounded to one decimal place for clarity.
Method RS Med AD Sci Fin
Zero-shot 32.3 28.3 15.6 35.6 62.6
Oracle 81.1 65.8 54.2 56.9 91.1
LoRA-FT Hu et al. (2022) 69.7 41.6 25.4 40.9 87.5
O-LoRA Wang et al. (2023) 74.6 44.4 30.0 41.5 87.2
MoELoRA Chen et al. (2024b) 77.5 41.9 27.6 40.1 86.8
CL-MoE Huai et al. (2025) 71.3 46.8 26.3 41.2 88.7
HiDe Guo et al. (2025a) 74.3 49.0 33.2 38.5 81.6
SEFE Chen et al. (2025) 77.3 50.4 37.2 40.9 86.8
DISCO Guo et al. (2025b) 76.0 45.2 43.8 42.3 89.0
ModalPrompt Zeng et al. (2024) 53.6 45.7 40.8 41.8 87.8
L2P Wang et al. (2022) 63.8 34.6 23.0 38.6 93.0
MR-LoRA Zhao et al. (2025) 80.9 65.3 54.1 56.7 91.1
Routing (Ours) 81.0 66.6 53.4 56.4 91.1
5.3 Standardized Evaluation on MLLM-CL Benchmark
To further validate our framework, we evaluate it against the MLLM-CL benchmark, which focuses on various domains such as Remote Sensing and Medical imaging. As summarized in Table 3, our routing mechanism effectively bridges the gap between zero-shot baselines and oracle.
The results highlight two primary strengths of our approach: First, by isolating domain-specific features into dedicated expert subspaces, our method achieves scores of 81.0 in Remote Sensing and 66.6 in Medical imaging, matching or slightly exceeding oracle performance. This indicates that modular specialization can be more effective than joint training, which often suffers from destructive gradient interference. Second, our approach maintains high efficiency by delivering competitive results—such as 91.1 in Finance and 56.4 in Science—using a single-pass inference. This offers a significant computational advantage over state-of-the-art baselines like MR-LoRA, which require expensive dual-inference passes to achieve similar performance.
Unlike traditional parameter-efficient continual learning methods, our framework does not require access to historical data or previous task gradients. This independence, combined with the ability to perform cross-expert utilization at inference time, confirms that token-level routing is a robust and scalable solution for multimodal continual learning.
@@ 5.3 Standardized Evaluation on MLLM-CL Benchmark
5.4 Ablation Studies
In this section, we conduct a series of ablation experiments to evaluate the robustness, scalability, and computational efficiency of our routing-based framework.
Scalability and Expert Interference.
We first evaluate the stability of the routing mechanism as the expert pool grows. A critical challenge in modular architectures is ensuring that the addition of new task-specific experts does not induce performance degradation.
We begin with a base configuration routing between two experts: SNLI and COCO. As indicated in Table 4, this setup yields substantial improvements on target tasks while maintaining near-baseline performance on unrelated tasks like Hateful Memes (HM) and MMIMDb.
The incremental addition of the HM module leads to a substantial gain in domain-specific performance (60.2 → 71.6), demonstrating the router’s capacity to correctly activate the relevant expert. Similarly, including the MMIMDb module significantly improves performance in that domain (46.6 → 63.2) while maintaining the established performance levels of prior tasks.
To assess resilience against irrelevant expertise, we introduced a LoRA module trained on ChartGemma Masry et al. (2025), a dataset unrelated to the core evaluation suite. As shown in Table 4, the inclusion of this unrelated expert does not degrade performance across existing tasks, reinforcing the framework’s scalability and its ability to ignore distractor modules.
Finally, we introduced an expert for XNLI Conneau et al. (2018), a text-only natural language inference task closely related to the multimodal SNLI. Interestingly, the addition of XNLI not only boosted performance on its own task but also improved scores on related image-text tasks (80.3 → 81.2). This suggests a positive transfer effect, where the routing mechanism effectively leverages cross-modal similarities to enhance reasoning across related domains.
Table 5: Computational efficiency comparison across methods. Training time measured in GPU hours on 32 A10 GPUs. Data requirements shown as relative factors, with the first task as baseline.
Method Training Time (hrs) Data Requirements
First Task All Tasks First Task All Tasks
Seq-FT 2.5 10.0
Seq-FT+ER 2.5 13.5 2.2×
MTL 2.5 28.7
Routing (Ours) 2.5 10.6
Computational Efficiency Analysis.
We analyze the computational advantages of our routing-based approach compared to standard continual learning methods and existing routing baselines. Table 5 summarizes training time and data requirements for InternVL2-2B across a four-task sequence.
As shown in Table 5, our framework is significantly more efficient than MTL and Experience Replay (ER). Beyond these standard baselines, our method offers two distinct advantages over recent modular approaches that utilize external router modules.
First, our method maintains data independence. Many existing routing frameworks require saving a subset of data Zhao et al. (2025) (e.g., 50 samples per task) from all past and current tasks to train a centralized router LoRA module. This creates a persistent dependency on historical data, violating the strict O(1) data scaling requirement of true continual learning. In contrast, our routing vectors are trained in isolation for each task, requiring zero access to prior data.
Second, our framework utilizes single-pass inference where based on input embeddings routing vector determine which LoRA module to be used. Methods such as MR-LoRA Zhao et al. (2025) methods often necessitate a "double inference" call: an initial pass to determine the appropriate expert via a router module, followed by a second pass to generate the response using the selected expert.
5.5 Routing Enables Cross-modal Transfer
Building upon the positive transfer observed between the text-only XNLI and multimodal SNLI tasks in Table 4, we further investigate the effectiveness of routing in facilitating cross-modal synergy. We evaluate this capability on the multi-lingual math dataset MGSM Shi et al. (2022), specifically focusing on zero-shot reasoning performance.
As detailed in Table 6, our experimental setup utilizes two distinct LoRA modules: orca-math, which is specialized for text-based mathematical reasoning, and multi-math, which is trained on both image and text modalities. While the text-centric orca-math expert naturally outperforms the multimodal variant on this specific task, the routing mechanism achieves a superior aggregate performance (0.82 vs. 0.78 for English).
Table 6: Cross-modal experiments on the MGSM dataset. Routing leverages modality-specific experts (Text-only vs. Vision + Text) to achieve superior performance, demonstrating successful knowledge transfer.
@@ 5.5 Routing Enables Cross-modal Transfer
orca-math (T) multi-math (V+T)
@@
Notably, the routed model often surpasses the performance of individual specialized experts, demonstrating that our framework does not merely select the most relevant module but dynamically synthesizes complementary information across modalities. By integrating the specialized reasoning of text-only experts with the alignment capabilities of multimodal experts, the routing mechanism effectively expands the model’s performance.
6 Discussion
While our routing approach shares architectural similarities with Mixture of Experts (MoE) systems like MoE-LLaVA Lin et al. (2024), it fundamentally differs in both purpose and implementation. Traditional MoE approaches improve model capacity through joint training of specialized experts within a unified architecture, requiring simultaneous access to all training data and joint optimization of experts and gating networks.
@@ 6 Discussion
Our method diverges from this paradigm in three key aspects. First, our experts (LoRA modules) are trained independently and sequentially, with each expert specializing on a single task without access to other tasks’ data. Second, we explicitly decouple expert learning from routing by first training task-specific LoRA modules, then learning lightweight routing vectors while keeping LoRA parameters frozen, unlike traditional MoEs where expert structure and routing are jointly optimized. Third, our approach requires only new task data, eliminating the need to retain or synthesize previous task data.
7 Conclusion
In this work, we address the challenge of catastrophic forgetting in MLLMs by introducing a token-level routing framework that seamlessly integrates new task-specific experts while strictly preserving foundational knowledge. Our approach distinguishes itself from traditional multi-task learning and experience replay by maintaining fixed resource requirements per task and eliminating the dependency on historical data, thus overcoming the linear computational growth that typically limits large-scale continual learning. Through extensive evaluations, we demonstrate that routing achieves performance parity with the theoretical multi-task upper bound on specialized tasks while simultaneously protecting general-purpose reasoning capabilities. Notably, we uncover that this modular architecture facilitates cross-modal transfer, allowing unimodal expertise to bolster multimodal reasoning, and exhibits a positive scaling trend where larger models demonstrate superior resistance to forgetting. Unlike existing routing baselines that require multi-pass inference and data replay, our single-pass mechanism provides an efficient and scalable solution for real-world deployment. Future research will explore the extension of this modularity to support cross-lingual transfer in massively multilingual multimodal settings.
8 Limitations
While our routing-based framework offers a scalable solution for multimodal continual learning, some limitations merit further investigation. First, the current implementation necessitates the simultaneous loading of multiple LoRA adapters during inference to enable token-level dynamic composition. Although each adapter is parameter-efficient , the cumulative memory footprint scales linearly with the number of tasks, which may present deployment challenges on hardware with restricted VRAM. Future work could explore hot-swapping mechanisms to dynamically load and unload adapters from system RAM to GPU memory based on high-level task identifiers or predicted routing requirements, thereby maintaining a near-constant inference memory overhead. Second, our training strategy prioritizes stability by isolating experts, which inadvertently minimizes sequential transfer during the training phase. Because each LoRA module is trained independently on its respective task without access to prior adapters, the model cannot inherently leverage "forward transfer" to accelerate the acquisition of subsequent, related skills during the optimization process. While we observe positive transfer at inference time through token-level routing , the training process remains strictly decoupled. Further research is needed to determine if initializing new experts via knowledge distillation from existing ones could facilitate collaborative learning without reintroducing catastrophic forgetting.